OpenStudy (anonymous):

A research technician is asked to prepare 120 mL of a 2.5 % aqueous solution of a drug from a stock solution. The stock solution contains 60 g per L of the drug. Show all calculations and units. A) What percent concentration is the stock solution? (1) B) How much of the stock solution and distilled water should be used to achieve the desired dilution?

OpenStudy (australopithecus):

120mL of a 2.5% solution convert the what I assume is %mass volume (after dividing by 100 you will have units g/mL) concentration to the same units as the stock solution. 2.5%/100 = 0.025g/mL convert mL into litres 1L = 1000mL using a ratio so units cancel out like numbers would (1000mL/1L)* 0.025g/mL (Notice mL cancel out you are left with grams per litre) = 25g/L Remember stock contains 60g per L Again just use a simple ratio \[\frac{25g/L}{60g/L}*120mL = 50mL\] Notice the units cancel out on concentration and you are left with mL Hope this is helpful. If you need further explanation feel free to ask and thanks for the question

OpenStudy (australopithecus):

Oh to convert the stalk solution to percentage it is pretty easy you need to convert Liters to mL again same ratio as we used before but flipped so liters cancel out (60g/L)*(1L/1000mL) = 0.06g/mL Now just multiply by 100 to convert to %(mass/volume) 0.06g/mL(100) = 6%